OpenStudy (anonymous):

Put the equation in the form ax2+bx+c=0. x^2+8x=-3x^2+5

OpenStudy (anonymous):

x^2+8x=-3x^2+5 x^2 + 8 x - (-3 x^2 + 5) = 0 4 x^2+8 x-5=0

OpenStudy (anonymous):

put the like terms together x^2+3x^2+8x-5=0 4x^2+8x-5=0

OpenStudy (anonymous):

Thanks guys!